OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively. an = 7 • (-3)n + 1 an = 7 • 3n - 1 an = 7 • (-3)n - 1 an = 7 • 3n

OpenStudy (jdoe0001):

\(\begin{matrix} \pm 3^{\square}&\pm 3^{\square}&\pm 3^{\square}&\pm 3^{\square}&\pm 3^{\square}\\ \square&-21&\square&\square&567\\ \hline\\ 1&2&3&4&5 \end{matrix}\)

OpenStudy (jdoe0001):

so what do you think?

OpenStudy (anonymous):

idk

OpenStudy (jdoe0001):

let's try the 2nd term is -21 based on the given choices, we know the common ratio has something to do with "3" 3 x 7 = 21 -3 x 7 = -21

OpenStudy (anonymous):

ok that makes sense so now im left with two options an = 7 • (-3)n + 1 an = 7 • (-3)n - 1 how do i get the second part?

OpenStudy (jdoe0001):

-21 is the 2nd term \(\begin{matrix} -3^\color{red}{1}\\ 21\\ \color{red}{2} \end{matrix}\)

OpenStudy (jdoe0001):

let's see the 5th term 567 \(3^4 = 81\\ 81 \times 7 = 567\\ (-3)^4 = 81\\ 81 \times 7 = 567\\ \begin{matrix} -3^\color{red}{4}\\ 21\\ \color{red}{5th} \end{matrix}\)

OpenStudy (jdoe0001):

darn, a typo

OpenStudy (jdoe0001):

\(3^4 = 81 \implies 81 \times 7 = 567\\ (-3)^4 = 81\\ 81 \times 7 = 567\\ \begin{matrix} -3^\color{red}{4}\\ 567\\ \color{red}{5th} \end{matrix}\)

OpenStudy (jdoe0001):

\(\begin{matrix} (-3)^{\color{red}{0}}&(-3)^{\color{red}{1}}&(-3)^{\color{red}{2}}&(-3)^{\color{red}{3}}&(-3)^{\color{red}{4}}\\ 7&-21&\square&\square&567\\ \color{red}{1}st&\color{red}{2}nd&\color{red}{3}rd&\color{red}{4}th&\color{red}{5}th \end{matrix}\)

OpenStudy (jdoe0001):

so, what do you think?

OpenStudy (jdoe0001):

what do you think would be the 3rd term?

OpenStudy (anonymous):

63???

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\begin{matrix} 7(-3)^{\color{red}{0}}&7(-3)^{\color{red}{1}}&7(-3)^{\color{red}{2}}&7(-3)^{\color{red}{3}}&7(-3)^{\color{red}{4}}\\ 7&-21&63&\square&567\\ \color{red}{1}st&\color{red}{2}nd&\color{red}{3}rd&\color{red}{4}th&\color{red}{5}th \end{matrix}\) yeap

OpenStudy (jdoe0001):

-3*-3 = +9 * 7 = +63 so there, that's the equation :) \(\bf 7\times (-3)^{n-1}\)